OpenStudy (anonymous):

help? xl2l4l6l8 f(x)l5l10l15l20 B.write an equation of the line that represents the function.

OpenStudy (anonymous):

that's a grahph fyi

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what @henryli

OpenStudy (johnweldon1993):

Okay so a graph with points (2,5) (4,10) (6,15) (8,20) So...we have at least 2 points...so lets find the slope... Lets take (2,5) and (4,10) Let 2 = x1 4 = x2 5 = y1 and 10 = y2 \[\large slope = m = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

thank you @johnweldon1993

OpenStudy (johnweldon1993):

And then with that slope 'm' plug that and any 1 point (Xo,Yo) into the point slope formula \[\large y - y_0 = m(x - x_0)\]

OpenStudy (anonymous):

okay